# SongEditor: Adapting Zero-Shot Song Generation Language Model as a Multi-Task Editor

### v1

@@ Abstract
The emergence of novel generative modeling paradigms, particularly audio language models, has significantly advanced the field of song generation. Although state-of-the-art models are capable of synthesizing both vocals and accompaniment tracks up to several minutes long concurrently, research about partial adjustments or editing of existing songs is still underexplored, which allows for more flexible and effective production. In this paper, we present SongEditor, the first song editing paradigm that introduces the editing capabilities into language-modeling song generation approaches, facilitating both segment-wise and track-wise modifications. SongEditor offers the flexibility to adjust lyrics, vocals, and accompaniments, as well as synthesizing songs from scratch. The core components of SongEditor include a music tokenizer, an autoregressive language model, and a diffusion generator, enabling generating an entire section, masked lyrics, or even separated vocals and background music. Extensive experiments demonstrate that the proposed SongEditor achieves exceptional performance in end-to-end song editing, as evidenced by both objective and subjective metrics. Audio samples are available in https://cypress-yang.github.io/SongEditor˙demo/.
@@ Introduction
As shown in figure 1, in addition to generating songs from scratch, SongEditor can also regenerate specific periods and separate vocals or accompaniments. Worth noting is that the above two capabilities can be performed at the same time, allowing for the editing of even a short segment of vocals or accompaniment.
The main contributions of this paper are summarized as follows:
@@
  • SongEditor is among the first works designed not only to generate songs from scratch, but also to conduct sentence-level editing at specific segment (segment-wise) and complete vocal or accompaniment tracks when given the rest (track-wise).
@@
Moreover, some approaches have extensively explored for the task of speech editing. For instance, VoiceCraft (Peng et al. 2024) proposes a token rearrangement method for the autoregressive generation procedure in neural codec language models. Apart from this, diffusion models (Vyas et al. 2023; Wang et al. 2023b; Du et al. 2024a) naturally possess the infilling editing ability in a non-autoregressive manner. These approaches have achieved a high quality of speech resynthesis. However, since the speech utterances used in these studies are relatively short and simple, they cannot be directly employed for song editing.
@@
Conventional SVS systems primarily focus on incorporating phoneme-level melody controls into the generated acoustic representations. For instance, VISinger (Zhang et al. 2022a, b) adopts a methodology similar to VITS (Kim, Kong, and Son 2021), which learns the alignments between acoustic and textual units. TokSing (Wu et al. 2024b) discretizes raw audio using multiple self-supervised learning models and further enhances melody information during both the training and inference phases. Recent models (Liu et al. 2022; Hwang, Lee, and Lee 2025; He et al. 2023) have demonstrated promising and competitive performance for high-quality singing voice generation. HiddenSinger (Hwang, Lee, and Lee 2025), for instance, utilizes the latent diffusion model (Rombach et al. 2022) to synthesize regularized intermediate codec representations. Meanwhile, RMSSinger (He et al. 2023) proposes a DDPM-based method (Ho, Jain, and Abbeel 2020) that synthesizes voice using word-level coarse-grained musical scores, alleviating the heavy demand for manual annotations in real-world scenarios.
@@
Notably, due to the absence of related works in the field of song generation and the fact that most commercial applications do not expose technical details, before introducing the editing framework, we would first like to demonstrate the base system, SongLM, for our editing paradigm. Figure 2 illustrates the framework of SongLM, which accepts multiple sentences of lyric, and a 10-second acoustic prompt as inputs. The main components include a semantic tokenizer, a language model, and a diffusion-based generator. The semantic tokenizer first compresses audio waveforms (including acoustic prompt, contextual waveforms and separated tracks) into discrete semantic tokens. Then the decoder-only transformer-based language model generates the semantic tokens autoregressively. The final waveform can be reconstructed from the output sequence by the generator, which consists of a diffusion transformer and an acoustic vocoder.
The training procedure can be divided into two stages: first, we train the tokenizer and diffusion generator simultaneously; then we train the language model with tokens derived from the tokenizer. Both the tokenizer and the generator are frozen in the second stage.
@@
The original data pairs only contain plain text and waveform $\{(X_{i},W_{i})\}_{i=1}^{L}$, where $L$ is the number of lyrics, $X_{i}$ is the $i$-th sentence and $W_{i}$ is the correspoonding waveform. It is worth noting that both waveforms and lyrics are consecutive, so they can be concatenated in sequence and $W_{i}$ may be empty at non-vocal sections. However, solely relying on the text of lyrics is not enough for the complicated song generation task. In order to incorporate the structure information into the lyrics, a structure detector  (Kim and Nam 2023) is applied. The category of structure indicators is shown in Appendix A.
@@
We utilize the Latent Diffusion Model (LDM)(Rombach et al. 2022) as the generator, comprising a diffusion model, a variational autoencoder (VAE)(Kingma and Welling 2013), and a vocoder. We replace the conventional U-Net (Ronneberger, Fischer, and Brox 2015) backbone with DiT (Peebles and Xie 2023), an advanced transformer-based diffusion model. DiT conditions on semantic tokens and applies forward and reverse processes on the latent vectors produced by VAE (van den Oord, Vinyals, and kavukcuoglu 2017). The VAE then converts these latent vectors into a Mel spectrogram, which is subsequently transformed into a waveform by the HifiGAN (Kong, Kim, and Bae 2020) vocoder.
@@
A delayed pattern (Copet et al. 2023) is further applied to rearrange the $K$ RVQ tokens. Tokens of the $k$-th quantizer are moved $(k-1)$ timesteps backward and extra $(K-1)$ empty frames are appended to each segment in order to prevent overlap of quantizers. Then tokens at the same timestep are stacked together. Consequently, the final length of the language model input should be $(T+3K)$. Because the lyrics of contexts have been discarded, only the training loss of editing segment is calculated as $\mathbb{L}_{ce}=-\log\left(P_{\theta}^{[A,B]}\right)$, where
@@
To achieve more robust and smooth transitions, a score-based candidate selection (Jiang et al. 2023) is applied during inference. Specifically, we first generate a candidate $\hat{y}^{1}_{[A:B]}$ and resynthesize the last 3 seconds for $(N-1)$ times, thus obtaining a set of $N$ candidates $\hat{Y}_{[A:B]}=\{\hat{y}^{1}_{[A:B]},\hat{y}^{2}_{[A:B]},\cdots,\hat{y}^{N}_{[A:B]}\}$. Then we utilize each candidate as new prefix, continuously predicting the next $\lambda$ steps. The score of each candidate can be represented as the log-likelihood of $y_{(B:B+\lambda]}$. Candidate with the largest score will be selected, which can be formulated as:
@@
Furthermore, we explore adding vocal or accompaniment completion functions to the model. As shown in Figure 2, a source separation algorithm is first applied to the original audio. Then, SongEditor takes either separated track as auxiliary context to complete the other one. The provided track is passed through a gated multi-source encoder and incorporated into the language model. Details are introduced below.
@@
The Band-Split RNN (BS-RNN) (Yu et al. 2023; Luo and Yu 2023) is adopted as source separation module for track-wise editing. BSRNN splits the spectrogram into multiple subbands and performs bidirectional LSTM layers across both subbands and frames. Given that music typically has a higher sample rate and a broader frequency range, BS-RNN is exceptionally suitable for music source separation. We leverage BS-RNN to separate vocals from the mixture, leaving the remaining as accompaniments.
@@
The proposed models are evaluated in terms of both objective and subjective metrics. For objective evaluation, both Phoneme Error Rate (PER) and Fréchet Audio Distance (FAD) (Kilgour et al. 2019) are used. To calculate PER, the vocal is first extracted by BSRNN. Then the Whisper-large-v2 is employed for speech recognition. All punctuation and structure tokens in the lyrics have been removed. The FAD score is computed based on the internal representation from the last layer of MERT-95M, and the fadtk [^2] https://github.com/microsoft/fadtk toolkit is used for calculating statistics and comparison. For editing tasks, only the editing segments are taken into consideration.

### v2

@@ Abstract
The emergence of novel generative modeling paradigms, particularly audio language models, has significantly advanced the field of song generation. Although state-of-the-art models are capable of synthesizing both vocals and accompaniment tracks up to several minutes long concurrently, research about partial adjustments or editing of existing songs is still underexplored, which allows for more flexible and effective production. In this paper, we present SongEditor, the first song editing paradigm that introduces the editing capabilities into language-modeling song generation approaches, facilitating both segment-wise and track-wise modifications. SongEditor offers the flexibility to adjust lyrics, vocals, and accompaniments, as well as synthesizing songs from scratch. The core components of SongEditor include a music tokenizer, an autoregressive language model, and a diffusion generator, enabling generating an entire section, masked lyrics, or even separated vocals and background music. Extensive experiments demonstrate that the proposed SongEditor achieves exceptional performance in end-to-end song editing, as evidenced by both objective and subjective metrics.
Demohttps://cypress-yang.github.io/SongEditor˙demo/
@@ Introduction
As shown in figure 1, in addition to generating songs from scratch, SongEditor can also regenerate specific periods and separate vocals or accompaniments. Worth noting is that the above two capabilities can be performed at the same time, allowing for the editing of even a short segment of vocals or accompaniment. The main contributions of this paper are summarized as follows:
@@
  • SongEditor is among the first works designed not only to generate songs from scratch, but also to conduct sentence-level editing at the specific segment (segment-wise) and complete vocal or accompaniment tracks when given the rest (track-wise).
@@
Moreover, some approaches have extensively explored the task of speech editing. For instance, VoiceCraft (Peng et al. 2024) proposes a token rearrangement method for the autoregressive generation procedure in neural codec language models. Apart from this, diffusion models (Vyas et al. 2023; Wang et al. 2023b; Du et al. 2024a) naturally possess the infilling editing ability in a non-autoregressive manner. These approaches have achieved a high quality of speech resynthesis. However, since the speech utterances used in these studies are relatively short and simple, they cannot be directly employed for song editing.
@@
Conventional SVS systems primarily focus on incorporating phoneme-level melody controls into the generated acoustic representations. For instance, VISinger (Zhang et al. 2022a, b) adopts a methodology similar to VITS (Kim, Kong, and Son 2021), which learns the alignments between acoustic and textual units. TokSing (Wu et al. 2024b) discretizes raw audio using multiple self-supervised learning models and further enhances melody information during both the training and inference phases. Recent models (Liu et al. 2022; Hwang, Lee, and Lee 2025; He et al. 2023) have demonstrated promising performance for high-quality singing voice generation. HiddenSinger (Hwang, Lee, and Lee 2025), for instance, utilizes the latent diffusion model (Rombach et al. 2022) to synthesize regularized intermediate codec representations. Meanwhile, RMSSinger (He et al. 2023) proposes a DDPM-based method (Ho, Jain, and Abbeel 2020) that synthesizes voice using word-level coarse-grained musical scores, alleviating the heavy demand for manual annotations in real-world scenarios.
@@
Before introducing the editing framework, we would first like to demonstrate the base system, SongLM, for our editing paradigm. Figure 2 presents the architectural framework of SongLM, which accepts multiple lyric sentences and a 10-second acoustic prompt as inputs. The primary components of this framework include a semantic tokenizer, a language model, and a diffusion-based generator. The semantic tokenizer compresses audio waveforms including the acoustic prompt, context, and separated tracks, into discrete semantic tokens. Subsequently, the language model generates semantic tokens in an autoregressive manner. The final waveform is reconstructed from the output sequence by the diffusion generator.
The training process is divided into two distinct phases. Initially, the tokenizer and diffusion generator are trained concurrently. Subsequently, the language model is trained using tokens generated by the tokenizer. During the second phase, both the tokenizer and the diffusion generator remain in a frozen state.
@@
The original data pairs only contain plain text and waveform $\{(X_{i},W_{i})\}_{i=1}^{L}$, where $L$ is the number of lyrics, $X_{i}$ is the $i$-th sentence and $W_{i}$ is the corresponding waveform. It is worth noting that both waveforms and lyrics are consecutive, so they can be concatenated in sequence and $W_{i}$ may be empty at non-vocal sections. However, relying solely on the text of lyrics is not enough for the complicated song generation task. In order to incorporate the structure information into the lyrics, a structure detector  (Kim and Nam 2023) is applied. The category of structure indicators is shown in Appendix A.
@@
We utilize the Latent Diffusion Model (LDM)(Rombach et al. 2022) as the generator, comprising a diffusion model, a variational autoencoder (VAE)(Kingma and Welling 2013), and a vocoder. We replace the conventional U-Net (Ronneberger, Fischer, and Brox 2015) backbone with DiT (Peebles and Xie 2023), which conditions on semantic tokens and applies forward and reverse processes on the latent vectors produced by VAE (van den Oord, Vinyals, and kavukcuoglu 2017). The VAE then converts latent vectors into a Mel spectrogram, which is subsequently transformed into a waveform by the HifiGAN (Kong, Kim, and Bae 2020) vocoder.
@@
A delayed pattern (Copet et al. 2023) is further applied to rearrange the $K$ RVQ tokens. Tokens of the $k$-th quantizer are moved $(k-1)$ timesteps backward and extra $(K-1)$ empty frames are appended to each segment in order to prevent overlap of quantizers. Then tokens at the same timestep are stacked together. Consequently, the final length of the language model input should be $(T+3K)$. Because the lyrics of contexts have been discarded, only the training loss of the editing segment is calculated as $\mathbb{L}_{ce}=-\log\left(P_{\theta}^{[A,B]}\right)$, where
@@
To achieve better transitions, a score-based candidate selection (Jiang et al. 2023) is applied during inference. Specifically, we first generate an initial candidate $\hat{y}^{1}_{[A:B]}$ and resynthesize the last 3 seconds for $(N-1)$ times, resulting in a set of $N$ candidates $\hat{Y}_{[A:B]}=\{\hat{y}^{1}_{[A:B]},\hat{y}^{2}_{[A:B]},\cdots,\hat{y}^{N}_{[A:B]}\}$. Each candidate is used as a new prefix to predict the subsequent $\lambda$ frames. The score of each candidate can be represented as the log-likelihood of $y_{(B:B+\lambda]}$. The candidate with the highest score is ultimately selected, which can be formulated as:
@@
Furthermore, we explore the integration of vocal and accompaniment completion into the model. As depicted in Figure 2, SongEditor utilizes either the separated vocal or accompaniment track as auxiliary context to complete the other. The provided track is processed through a gated multi-source encoder and incorporated into the language model. Details of this process are introduced below.
@@
The Band-Split RNN (BS-RNN) (Yu et al. 2023; Luo and Yu 2023) is adopted as source separation module for track-wise editing. BS-RNN splits the spectrogram into multiple subbands and performs bidirectional LSTM layers across both subbands and frames. Given that music typically has a higher sample rate and a broader frequency range, BS-RNN is exceptionally suitable for music source separation. We leverage BS-RNN to separate vocals from the mixture, leaving the remaining as accompaniments.
@@
The proposed models are evaluated in terms of both objective and subjective metrics. For objective evaluation, both Phoneme Error Rate (PER) and Fréchet Audio Distance (FAD) (Kilgour et al. 2019) are employed. To calculate PER, the vocal is first extracted by BSRNN. Then the Whisper-large-v2 is utilized for speech recognition. All punctuation and structure tokens in the lyrics have been removed. The FAD score is computed based on the internal representation from the last layer of MERT-95M, and the fadtk [^2] https://github.com/microsoft/fadtk toolkit is used for calculating statistics and comparison. For editing tasks, only the editing segments are taken into consideration.